OpenStudy (anonymous):

find the equation of the line through (2,3) and (4,7)

OpenStudy (anonymous):

Use the slope equation: \[m=\frac{y_1-y_2}{x_1-x_2}\] \[m=\frac{7-3}{4-2}\] \[m=\frac{4}{2}=2\] Then plug it into the equation: \[y=mx+b\] You can use any of the point to substitute y and x. \[7=2(4)+b\] \[7-8=b=-1\] Therefore: \[y=2x-1\]

OpenStudy (anonymous):

how did you get 2x and can you explain "m=4/2=2 how did you get that =2

OpenStudy (anonymous):

@finaiized is correct!

OpenStudy (precal):

you can plot those two points and see that the equation does go through those 2 points

OpenStudy (anonymous):

The slope equation is \[y=mx+b\] I first substituted in an x and y from the points you gave me to find b. Once I found b to be -1, I generalized the equation by replacing the numbers I substituted with y and x. If you plug in any point in y and x, then you'll get a point on the line. As for the m=4/2, I subtracted points from each pair of points (Y1 from example, is the y coordinate of one of the points. Then I just subtracted Y1-Y2, and X1-X2, and got 4/2 which is the same as 2.

OpenStudy (precal):

I agree with finaiized

OpenStudy (anonymous):

make sense, thanks so much...i understand